# references.bib

### v2

@@ 5 $\mathllap{\text{{target\_nodes}}}$leftarrow batched_edge_index[1]
3 Methods
In this section, we provide a high-level overview of our architecture and describe the masked attention module that allows for information propagation across edges as primitives. An alternative implementation for the more traditional node-based propagation is also presented. The section starts with a formal definition of the masked self-attention mechanism and then proceeds to describe our end-to-end attention-based approach for learning on graphs, involving the encoder and pooling components (illustrated in Figure 3).
3.1 Masked Attention Modules
We first introduce the basic notation and then give a formal description of the masked attention mechanism with a focus on the connectivity pattern specific to graphs. Following this, we provide algorithms for an efficient implementation of the masking operators using native tensor operations.
As in most graph learning settings, a graph is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{E})$ where $\mathcal{N}$ represents the set of nodes (vertices), $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is the set of edges, and $N_{n}=|\mathcal{N}|$, $N_{e}=|\mathcal{E}|$. Nodes are associated with feature vectors $\mathbf{n}_{i}$ of dimension $d_{n}$ for all nodes $i\in\mathcal{N}$, and $d_{e}$-dimensional edge features $\mathbf{e}_{ij}$ for all edges $e\in\mathcal{E}$. The node features are collected as rows in a matrix $\mathbf{N}\in\mathbb{R}^{N_{n}\times d_{n}}$, and similarly for edge features into $\mathbf{E}\in\mathbb{R}^{N_{e}\times d_{e}}$. The graph connectivity information can be represented as an adjacency matrix $\mathbf{A}$, where $\mathbf{A}_{ij}=1$ if $(i,j)\in\mathcal{E}$ and $\mathbf{A}_{ij}=0$ otherwise. The edge list (edge index) representation is equivalent but more common in practice.
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are projections of the edge representations to queries, keys, and values, respectively. This is a minor extension of the standard scaled dot product attention via the additive mask $\mathbf{M}$ that can censor any of the pairwise attention scores, and $d_{k}$ is the key dimension. The generalisation to masked multihead attention follows the same steps as in the original transformer [NIPS2017_3f5ee243]. Below and in the specification of the overall architecture, we refer to this function as $\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{M})$.
Masked self-attention for graphs can be seen as graph-structured attention, and an instance of a generalized attention mechanism with custom attention patterns [shirzad2023exphormer]. More specifically, in ESA the attention pattern is given by an edge adjacency matrix rather than allowing for interactions between all set items. Crucially, the edge adjacency matrix can be efficiently computed both for a single graph and for batched graphs using exclusively tensor operations. The case for a single graph is covered in the left side of Section 2 through the edge_adjacency function. The first 3 lines of the function correspond to getting the number of edges, then separating the source and target nodes from the edge adjacency list (also called edge index), which is equivalent to the standard adjacency matrix of a graph. The source and target node tensors each have a dimension equal to the number of edges ($N_{e}$). Lines 7-9 add an additional dimension and efficiently repeat (‘expand’) the existing tensors to shape $(N_{e},N_{e})$ without allocating new memory. Using the transpose, line 11 checks if the source nodes of any two edges are the same, and the same for target nodes on line 12. On line 13, cross connections are checked, where the source node of an edge is the target node of another, and vice versa. The operations of lines 11-14 result in boolean matrices of shape $(N_{e},N_{e})$ which are summed for the final returned edge adjacency matrix. The right panel of Section 2 depicts the case where the input graph represents a batch of smaller graphs. This requires an additional batch mapping tensor that maps each node in the batched graph to its original graph, and carefully manipulating the indices to create the final edge adjacency mask of shape $(B,L,L)$, where $L$ is the maximum number of edges in the batch.
Since in ESA attention is computed over edges, we chose to separate source and target node features for each edge, similarly to lines 4-5 of Section 2, and concatenate them to the edge features:
Having defined the mask generation process and the masked multihead attention function, we next define the modular blocks of ESA, starting with the Masked Self Attention Block (MAB):
$\mathbf{\overline{X}}$ $=\text{LayerNorm}(\mathbf{X})$ (5)
where the mask $\mathbf{M}$ is computed as in Section 2 and has shape $B\times L\times L$, with $B$ the batch size, and MLP is a multi-layer perceptron. A Self Attention Block (SAB) can be formally defined as:
$\text{SAB}(\mathbf{X})=\text{MAB}(\mathbf{X},\mathbf{0})$ (6)
1 from torch_geometric import unbatch_edge_index
2 function node_mask(batched_edge_index, batch_map, B, M)
       4 # batch_map maps nodes to graphs
       5 # B, M are the batch, respectively mask size
       6 mask $\leftarrow$ torch.full(size=(B, M, M), fill=False)
       9 edge_index $\leftarrow$ torch.cat(edge_index_list, dim=1)
[…]
Algorithm 1 Node masking in the PyTorch Geometric framework.
3.2 ESA Architecture
Our architecture consists of two components: i) an encoder that interleaves masked and self-attention blocks to learn an effective representation of edges, and ii) a pooling block based on multi-head attention, inspired by the decoder component from the set transformer architecture [lee2019set]. The latter comes naturally when one considers graphs as sets of edges (or nodes), as done here. ESA is a purely attention-based architecture that leverages the scaled dot product mechanism proposed by Vaswani et al. [NIPS2017_3f5ee243] through masked- and standard self-attention blocks (MABs and SABs) and a pooling by multi-head attention (PMA) block.
Here, ‘AB’ refers to an attention block that can be instantiated as an MAB or SAB.
where $\mathbf{S}_{k}$ is a tensor of $k$ learnable seed vectors that are randomly initialised and $\text{SAB}^{p}(\cdot)$ is the application of $p$ SABs. Technically, it suffices to set $k=1$ to output a single representation for the entire graph. However, we have empirically found it beneficial to set it to a small value, such as $k=32$. Moreover, this change allows self attention (SABs) to further process the $k$ resulting representations, which can be simply summed or averaged due to the small $k$. Contrary to classical readouts that aggregate directly over set items (i.e., nodes), pooling by multi-head attention performs the final aggregation over the embeddings of learnt seed vectors $S_{k}$. While tasks involving node-level predictions require only the Encoder component, the predictive tasks involving graph-level representations require all the modules, both the encoder and pooling by multi-head attention. The architecture in the latter setting is formally given by
$\mathbf{Z}_{\text{out}}=\text{PMA}_{k,p}(\text{Encoder}(\mathbf{X},\mathbf{M})+\mathbf{X})$
As optimal configurations are task specific, we do not explicitly fix all the architectural details. For example, it is possible to select between layer and batch normalisation, a pre-LN or post-LN architecture [10.5555/3524938.3525913], or standard and gated MLPs, along with GLU variants, e.g., SwiGLU [DBLP:journals/corr/abs-2002-05202].
3.3 Time and Memory Scaling
ESA is enabled by recent advances in efficient and exact attention, such as memory-efficient attention and Flash attention [DBLP:journals/corr/abs-2112-05682, dao2022flashattention, dao2023flashattention2, shah2024flashattention3fastaccurateattention]. Flash attention does not yet support masking, but memory-efficient implementations with arbitrary masking capabilities exist in both PyTorch and xFormers [NEURIPS2019_9015, xFormers2022]. Theoretically, the memory complexity of memory-efficient approaches is $O(\sqrt{n})$, where $n$ is the sequence/set size, with the same time complexity as standard attention. Since all operations in ESA except the cross-attention in PMA are based on self-attention, our method benefits directly from all of these advances. Moreover, our cross-attention is performed between the full set of size $n$ and a small set of $k\leq 32$ learnable seeds, such that the complexity of the operation is $O(kn)$ even for standard attention. Flash attention, which has linear memory complexity even for self-attention and is up to ×$\times 25$ times faster than standard attention, supports cross-attention and we use it for further uplifts. However, these implementations are not optimised for graph learning and we have found that the biggest bottleneck is not the attention computation, but rather storing the dense edge adjacency mask, which requires memory quadratic in the number of edges. An evident, but not yet available, optimisation could amount to storing the masks in a sparse tensor format. Alternatively, even a single boolean requires an entire byte of storage in PyTorch, increasing the theoretical memory usage by 8 times. Further limitations and possible optimisations are discussed in SI 5. Despite some of these limitations, we have successfully trained ESA models for graphs with up to approximately 30,000 edges (e.g., dd in Table 6).
4 Results
@@
We perform a comprehensive evaluation of ESA on 70 different tasks, including domains such as molecular property prediction, vision graphs, and social networks, as well as different aspects of representation learning on graphs, ranging from node-level tasks with homophily and heterophily graph types to modelling long range dependencies, shortest paths, and 3D atomic systems. We quantify the performance of our approach relative to 6 GNN baselines: GCN, GAT, GATv2, PNA, GIN, and DropGIN (more expressive than $1$-WL), and 3 graph transformer baselines: Graphormer, TokenGT, and GraphGPS. All the details on hyper-parameter tuning, rationale for the selected metrics, and selection of baselines can be found in Sections SI 7.1, SI 7.2 and SI 7.3. In the remainder of the section, we summarize our findings across molecular learning, mixed graph-level tasks, node-level tasks, and ablations on the interleaving operator along with insights on time and memory scaling.
4.1 Molecular Learning
@@
As learning on molecules has emerged as one of the most successful applications of graph learning, we present an in-depth evaluation including quantum mechanics, molecular docking, and various physical chemistry and biophysics benchmarks, as well as an exploration of learning on 3D atomic systems, transfer learning, and learning on large molecules of therapeutic relevance (peptides).
@@
Table 1: The table reports root mean squared error on qm9 (RMSE is the standard metric for quantum mechanics) and $\text{R}^{2}$ for dockstring (dock) and MoleculeNet (mn), presented as mean $\pm$ standard deviation over $5$ runs. The mean absolute error (MAE) is reported for pcqm4mv2 (MAE is the standard metric for this task) over a single run due to the size of the dataset. The lowest MAEs and RMSEs, and the highest $\text{R}^{2}$ values are highlighted in bold. oom denotes out-of-memory errors. A complete table, including GCN and GIN, is provided in Table 1.
@@
Table 2: The table reports mean absolute error (MAE) on the zinc dataset, presented as mean $\pm$ standard deviation over $5$ runs. The best/lowest value is highlighted in bold. A complete table, including GCN, GIN, and DropGIN is provided in Table 2.
@@
QM9. We report results for all $19$ qm9 [Ramakrishnan2014] targets in Table 1, with GCN and GIN separately in Table 1 due to space restrictions. We observe that on $15$ out of $19$ properties, ESA is the best performing model. The exceptions are the frontier orbital energies (HOMO and LUMO energy, HOMO-LUMO gap) and the dipole moment ($\mu$), where PNA is slightly ahead of it. Other graph transformers are competitive relative to GNNs on many properties but vary in performance across tasks.
@@
DOCKSTRING. dockstring [doi:10.1021/acs.jcim.1c01334] is a recent drug discovery data collection consisting of molecular docking scores for 260,155 small molecules and 5 high-quality targets from different protein families that were selected as a regression benchmark, with different levels of difficulty: parp1 (enzyme, easy), f2 (protease, easy to medium), kit (kinase, medium), esr2 (nuclear receptor, hard), and pgr (nuclear receptor, hard). We report results for the $5$ targets in Table 1 (and Table 1) and observe that ESA is the best performing method on 4 out of 5 tasks, with PNA slightly ahead on the medium-difficulty kit. TokenGT and GraphGPS do not generally match ESA or even PNA. Molecular docking scores also depend heavily on the 3D geometry, as discussed in the original paper [doi:10.1021/acs.jcim.1c01334], posing a difficult challenge for all methods. Interestingly, not only ESA but all tuned GNN baselines outperform the strongest method in the original manuscript (Attentive FP, a GNN based on attention [doi:10.1021/acs.jmedchem.9b00959]) despite using 20,000 less training molecules (which we leave out as a validation set). This illustrates the importance of evaluation relative to baselines with tuned hyper-parameters.
@@
MoleculeNet and NCI. We report results for a varied selection of three regression and three classification benchmarks from MoleculeNet, as well as two benchmarks from the National Cancer Institute (NCI) consisting of compounds screened for anti-cancer activity (Tables 1 and 3, and Tables 1 and 4). We also report the accuracy in Table 5. Except hiv, these datasets pose a challenge to graph transformers due to their small size ($<$5,000 compounds). With the exception of the Lipophilicity (lipo) dataset from MoleculeNet, we observe that ESA is the preferred method, and often by a significant margin, for example on bbbp and bace, despite their small size (2,039, respectively 1,513 total compounds before splitting). On the other hand, Graphormer and TokenGT perform poorly, possibly due to the small-sample nature of these tasks. GraphGPS is closer to the top performers but still underperforms compared to GAT(v2) and PNA. These results also show the importance of appropriate evaluation metrics, as the accuracy on hiv for all methods is above $97\%$ (Table 5), but the MCC (Table 3) is significantly lower.
@@
PCQM4MV2. pcqm4mv2 is a quantum chemistry benchmark introduced as a competition through the Open Graph Benchmark Large-Scale Challenge (OGB-LSC) project [hu2021ogblsc]. It consists of 3,378,606 training molecules with the goal of predicting the DFT-calculated HOMO-LUMO energy gap from 2D molecular graphs. It has been widely used in the literature, especially to evaluate graph transformers [ying2021do, kim2022pure, choi2024topologyinformedgraphtransformer]. Since the test splits are not public, we use the available validation set (73,545 molecules) as a test set, as is often done in the literature, and report results on it after training for 400 epochs. We report results from a single run, as it is common in the field due to the large size of the dataset [ying2021do, kim2022pure, choi2024topologyinformedgraphtransformer]. For the same reason, we do not run our baselines and instead chose to focus on the state-of-the-art results published on the official leaderboard [^1] https://ogb.stanford.edu/docs/lsc/leaderboards/#pcqm4mv2. At the time of writing (November 2024), the best performing model achieves a validation set MAE of 0.0671 [hussain2024triplet]. Here, ESA achieves a MAE of 0.0235, which is almost 3 times lower. It it worth noting that the top 3 methods for this dataset are all bespoke architectures designed for molecular learning, for example Uni-Mol+ [Lu2024], Transformer-M [luo2023one], and TGT [hussain2024triplet]. In contrast, ESA is general purpose (does not use any information or technique specifically for molecular learning), uses only the 2D input graph, and does not use any positional or structural encodings.
Table 3: The table reports Matthews correlation coefficient (MCC) for graph-level molecular classification tasks from MoleculeNet and National Cancer Institute (nci), presented as mean $\pm$ standard deviation over 5 different runs. oom denotes out-of-memory errors. The highest mean values are highlighted in bold. A complete table, including GCN and GIN, is provided in Table 4 with MCC as metric, and in Table 5 with accuracy.
Data ($\uparrow$) DropGIN GAT GATv2 PNA Graphormer TokenGT GPS ESA
Table 4: The table reports mean absolute error (MAE) and average precision (AP) for two long-range molecular benchmarks involving peptides. The molecular graph of a peptide is much larger than that of a small drug-like molecule and this makes the tasks well-suited for long-range benchmarking. All the results except the ones for ESA and TokenGT are extracted from [tonshoff2023where]. The number of layers for pept-struct, respectively pept-func is given as (/)($\cdot$/$\cdot$)( ⋅ / ⋅ ).
Dataset GCN (6/6) GIN (10/8) GPS (8/6) TokenGT (10/10) ESA (3/4)
pept-fn (AP $\uparrow$) $0.686$$\pm$ $0.005$ $0.6621$$\pm$ $0.0067$ $0.6534$$\pm$ $0.0091$ $0.6263465732336044$ $\pm$ $0.011736145935986075$ $0.686267567$$\pm$ $0.004352919$
@@
Table 5: A summary of the transfer learning performance on qm9 for HOMO and LUMO properties, presented as mean $\pm$ standard deviation over 5 different runs. The metric is the root mean squared error (RMSE). All the models use the 3D atomic coordinates and atom types as inputs and no other node or edge features. ‘Strat.’ stands for strategy and specifies the type of learning: GW only (no transfer learning), inductive, or transductive. The lowest values are highlighted in bold. A complete table, including GCN and GIN, is provided in Table 3.
@@
ZINC. We report results on the full zinc dataset with 250,000 compounds (Table 2), which is commonly used for generative purposes [pmlr-v80-jin18a, Kondratyev2023]. This is one of the only benchmarks where the graph transformer baselines (Graphormer, TokenGT, GraphGPS) convincingly outperformed strong GNN baselines. ESA, without positional encodings, slightly underperforms compared to GraphGPS, which uses random-walk structural encodings (RWSE). This type of encoding is known to be beneficial in molecular tasks, and especially for zinc [rampasek2022recipe, 10.5555/3648699.3648742]. Thus, we also evaluated an ESA + RWSE model, which increased relative performance by almost 40%. While there is no leaderboard available for the full version of zinc, recently Ma et al. [10.5555/3618408.3619379] evaluated their own method (GRIT) against 11 other baselines, including higher-order GNNs, with the best reported MAE being $\mathbf{0.023}\pm\mathbf{0.001}$. This shows that ESA can already almost match state-of-the-art models without structural encodings, and significantly improves upon this result when augmented.
@@
Long-range peptide tasks. Graph learning with transformers has traditionally been evaluated on long-range graph benchmarks (LRGB) [dwivedi2022long]. However, it was recently shown that simple graph neural networks outperform most attention-based methods [tonshoff2023where]. We selected two long-range benchmarks involving peptide property prediction: peptides-struct and peptides-func. From the LRGB collection, these two stand out due to having the longest average shortest path ($20.89\pm 9.79$ versus $10.74\pm 0.51$ for the next) and the largest average diameter ($56.99\pm 28.72$ versus $27.62\pm 2.13$). We report ESA results against tuned GNNs and GraphGPS models from [dwivedi2022long], as well as our own optimised TokenGT model (Table 4). Despite using only half the number of layers as other methods or less, ESA outperformed these baselines and matched the second model on the pept-struct leaderboard, and is within the top $5$ for pept-func (as of November 2024).
@@
Learning on 3D atomic systems. We adapt a subset from the Open Catalyst Project (OCP) [zitnick2020introduction, doi:10.1021/acscatal.0c04525] for evaluation, with the full steps in SI 9, including deriving edges from atom positions based on a cutoff, pre-processing specific to crystal systems, and encoding atomic distances using Gaussian basis functions. As a prototype, we compare NSA (MAE of $\mathbf{0.799}\pm\mathbf{0.008}$) against Graphormer ($0.839\pm 0.005$), one of the best models that have been used for the Open Catalyst Project [shi2023benchmarking]. These encouraging results on a subset of OCP motivated us to further study modelling 3D atomic system through the lens of transfer learning.
@@
Transfer learning on frontier orbital energies. We follow the recipe recently outlined for drug discovery and quantum mechanics by [Buterez_Janet_Kiddle_Oglic_Liò_2024] and leverage a recent, refined version of the qm9 HOMO and LUMO energies [Fediai2023] that provides alternative DFT calculations and new calculations at the more accurate GW level of theory. As outlined by [Buterez_Janet_Kiddle_Oglic_Liò_2024], transfer learning can occur transductively or inductively. The transfer learning scenario is important and is detailed in SI 4. We use a 25K/5K/10K train/validation/test split for the high-fidelity GW data, and we train separate low-fidelity DFT models on the entire dataset (transductive) or with the high-fidelity test set molecules removed (inductive). Since the HOMO and LUMO energies depend to a large extent on the molecular geometry, we use the 3D-aware version of ESA from the previous section, and adapt all of our baselines to the 3D setup. Our results are reported in Table 5. Without transfer learning (strategy ‘GW’ in Table 5), ESA and PNA are almost evenly matched, which is already an improvement since PNA was better for frontier orbital energies without 3D structures (Table 1), while the graph transformers perform poorly. Employing transfer learning all the methods improve significantly, but ESA outperforms all baselines for both HOMO and LUMO, in both transductive and inductive tasks.
4.2 Mixed Graph-level Tasks
@@
Other than molecular learning, we examine a suite of graph-level benchmarks from various domains, including computer vision, bioinformatics, synthetic graphs, and social graphs. Our results are reported in Tables 6 and 7, where ESA generally outperforms all the baselines. In the context of state-of-the-art models from online leaderboards, of particular note are the accuracy results (provided in Table 7) on the vision datasets, where on mnist ESA matches the best performing model [chen2024learninglongrangedependencies], and is in the top 5 on cifar10 at the time of submission (November 2024), without positional or structural encodings or other helper mechanisms. Similarly, on the malnettiny dataset of function call graphs popularised by GraphGPS [rampasek2022recipe], we achieve the highest recorded accuracy.
@@
Table 6: The table reports Matthews correlation coefficient (MCC) for graph-level classification tasks from various domains, presented as mean $\pm$ standard deviation over 5 different runs. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g., node/edge features are not integers, which are required for Graphormer and TokenGT). The poor performance of GraphGPS on malnettiny in Table 6 can be explained by our use of one-hot degrees as node features for datasets lacking pre-computed features, while GraphGPS originally used the more informative local degree profile [DBLP:journals/corr/abs-1811-03508]. The highest mean values are highlighted in bold. A complete table, including GCN and GIN, is provided in Table 6 with MCC as metric, and in Table 7 with accuracy.
@@
Table 7: The table reports Matthews correlation coefficient (MCC) for $11$ node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated ‘infected’ Erdős–Rényi (ER) graphs; details are provided in SI 8). For squirrel and chameleon, we used the filtered datasets introduced by Platonov et al. [platonov2023a] to fix existing data leaks. The highest mean values are highlighted in bold. Additional heterophily results are provided in Table 10. A complete table, including GIN and DropGIN, is provided in Table 8 with MCC as the performance metric, and in Table 9 with accuracy.
@@
Table 8: Example of multiple ESA configurations and their impact on performance via three different kinds of benchmarks: a graph-level molecular task (bbbp), a graph-level vision task (mnist), and a node-level heterophilous task (chameleon). In the model configurations, M denotes a MAB, S a SAB before the PMA module and S afterwards, and P is the PMA module. The performance metric is the Matthews correlation coefficient (MCC).

### v3

@@ 5 $\mathllap{\text{{target\_nodes}}}$leftarrow batched_edge_index[1]
Results
@@
We perform a comprehensive evaluation of ESA on 70 different tasks, including domains such as molecular property prediction, vision graphs, and social networks, as well as different aspects of representation learning on graphs, ranging from node-level tasks with homophily and heterophily graph types to modelling long range dependencies, shortest paths, and 3D atomic systems. We quantify the performance of our approach relative to 6 GNN baselines: GCN, GAT, GATv2, PNA, GIN, and DropGIN (more expressive than $1$-WL), and 3 graph transformer baselines: Graphormer, TokenGT, and GraphGPS. All the details on hyperparameter tuning, rationale for the selected metrics, and selection of baselines can be found in Sections SI 9.1, SI 9.2 and SI 9.3. In the remainder of the section, we summarize our findings across molecular learning, mixed graph-level tasks, node-level tasks, ablations on the interleaving operator along with insights on time and memory scaling, and explainability enabled by the learnt attention weights.
Molecular Learning
@@
As learning on molecules has emerged as one of the most successful applications of graph learning, we present an in-depth evaluation that includes quantum mechanics, molecular docking, and various physical chemistry and biophysics benchmarks, as well as an exploration of learning on 3D atomic systems, transfer learning, and learning on large molecules of therapeutic relevance (peptides).
@@
Table 1: The table reports the root mean squared error (RMSE) on qm9, the mean absolute error (MAE) for zinc, and $\text{R}^{2}$ for dockstring (dock) and MoleculeNet (mn), presented as mean $\pm$ standard deviation over $5$ runs. The MAE is reported for pcqm4mv2 (it is the standard metric for this task) over a single run due to the size of the dataset. The lowest MAEs and RMSEs, the highest $\text{R}^{2}$ values, and table headers are highlighted in bold. oom denotes out-of-memory errors. A complete table, including GCN and GIN, is provided in Table 1.
@@
Mean absolute error (MAE) on the full zinc dataset, including an ESA model augmented with positional encodings (PE). A complete table, including GCN, GIN, and DropGIN is provided in Table 3.
@@
QM9. We report results for all $19$ qm9 [Ramakrishnan2014] targets in Table 1, with GCN and GIN separately in Table 1 due to space restrictions. We observe that on $15$ out of $19$ properties, ESA is the best-performing model. The exceptions are the frontier orbital energies (HOMO and LUMO energy, HOMO-LUMO gap) and the dipole moment ($\mu$), where PNA is slightly ahead of it. Other graph transformers are competitive relative to GNNs on many properties, but vary in performance across tasks.
@@
DOCKSTRING. dockstring [doi:10.1021/acs.jcim.1c01334] is a recent drug discovery data collection consisting of molecular docking scores for 260,155 small molecules and 5 high-quality targets from different protein families that were selected as a regression benchmark, with different levels of difficulty: parp1 (enzyme, easy), f2 (protease, easy to medium), kit (kinase, medium), esr2 (nuclear receptor, hard), and pgr (nuclear receptor, hard). We report results for the 5 targets in Table 1 (and Table 1) and observe that ESA is the best performing method on 4 out of 5 tasks, with PNA slightly ahead on the medium-difficulty kit. TokenGT and GraphGPS generally do not match ESA or even PNA. Molecular docking scores also depend heavily on 3D geometry, as discussed in the original paper [doi:10.1021/acs.jcim.1c01334], posing a difficult challenge for all methods. Interestingly, not only ESA but all tuned GNN baselines outperform the strongest method in the original manuscript (Attentive FP, a GNN based on attention [doi:10.1021/acs.jmedchem.9b00959]) despite using 20,000 less training molecules (which we leave out as a validation set). This illustrates the importance of evaluating relative to baselines with tuned hyperparameters.
@@
MoleculeNet and NCI. We report results for a varied selection of three regression and three classification benchmarks from MoleculeNet, as well as two benchmarks from the National Cancer Institute (NCI) consisting of compounds screened for anti-cancer activity (Tables 1 and 3, and Tables 1 and 5). We also report the accuracy in Table 6. Apart from hiv, these datasets pose a challenge to graph transformers due to their small size ($<$5,000 compounds). With the exception of the Lipophilicity (lipo) dataset from MoleculeNet, we observe that ESA is the preferred method, and often by a significant margin, for example on bbbp and bace, despite their small size (2,039, respectively 1,513 total compounds before splitting). On the other hand, Graphormer and TokenGT perform poorly, possibly due to the small-sample nature of these tasks. GraphGPS is closer to the top performers, but still underperforms compared to GAT(v2) and PNA. These results also show the importance of appropriate evaluation metrics, as the accuracy on hiv for all methods is above $97\%$ (Table 6), but the MCC (Table 3) is significantly lower.
@@
PCQM4MV2. pcqm4mv2 is a quantum chemistry benchmark introduced as a competition through the Open Graph Benchmark Large-Scale Challenge (OGB-LSC) project [hu2021ogblsc]. It consists of 3,378,606 training molecules with the goal of predicting the DFT-calculated HOMO-LUMO energy gap from 2D molecular graphs. It has been widely used in the literature, especially to evaluate graph transformers [ying2021do, kim2022pure, pmlr-v251-choi24a]. Since the test splits are not public, we use the available validation set (73,545 molecules) as a test set, as is often done in the literature, and report results on it after training for 400 epochs. We report results from a single run, as it is common in the field due to the large size of the dataset [ying2021do, kim2022pure, pmlr-v251-choi24a]. For the same reason, we do not run our baselines and instead chose to focus on the state-of-the-art results published on the official leaderboard. At the time of writing (March 2025), the best performing model achieved a validation set MAE of 0.0671 [hussain2024triplet]. Here, ESA achieves a value of 0.0235, which is almost 3 times lower. It is worth noting that the top 3 methods for this dataset are all bespoke architectures designed for molecular learning, for example Uni-Mol+ [Lu2024], Transformer-M [luo2023one], and TGT [hussain2024triplet]. In contrast, ESA is general purpose (does not use any information or technique specifically for molecular learning), uses only the 2D input graph, and does not use any positional or structural encodings.
@@
Table 2: A summary of the transfer learning performance on qm9 for HOMO and LUMO properties, presented as mean $\pm$ standard deviation over 5 different runs. The metric is the root mean squared error (RMSE). All models use 3D atomic coordinates and atom types as inputs and no other node or edge features. ‘Strat.’ stands for strategy and specifies the type of learning: GW only (no transfer learning), inductive, or transductive. The lowest values and table headers are highlighted in bold. A complete table, including GCN and GIN, is provided in Table 4.
@@
ZINC. We report results on the full zinc dataset with 250,000 compounds (Table 1), which is commonly used for generative purposes [pmlr-v80-jin18a, Kondratyev2023]. This is one of the only benchmarks where the graph transformer baselines (Graphormer, TokenGT, GraphGPS) convincingly outperformed strong GNN baselines. ESA, without positional encodings, slightly underperforms compared to GraphGPS, which uses random-walk structural encodings (RWSE). This type of encoding is known to be beneficial in molecular tasks, and especially for zinc [rampasek2022recipe, 10.5555/3648699.3648742, grötschla2024benchmarkingpositionalencodingsgnns]. Thus, we also evaluated an ESA + RWSE model, which increased relative performance by almost 45%. Recently, Ma et al. [10.5555/3618408.3619379] evaluated their own method (GRIT) against 11 other baselines, including higher-order GNNs, with the best reported MAE being $0.023\pm 0.001$. This shows that ESA can already almost match state-of-the-art models without structural encodings, and significantly improves upon this result when augmented.
@@
Long-range peptide tasks. Graph learning with transformers has traditionally been evaluated on long-range graph benchmarks (LRGB) [dwivedi2022long]. However, it was recently shown that simple GNNs outperform most attention-based methods [tonshoff2023where]. We selected two long-range benchmarks that involve the prediction of peptide properties: peptides-struct and peptides-func. From the LRGB collection, these two stand out because they have the longest average shortest path ($20.89\pm 9.79$ versus $10.74\pm 0.51$ for the next) and the largest average diameter ($56.99\pm 28.72$ versus $27.62\pm 2.13$). We report ESA results against tuned GNNs and GraphGPS models from [dwivedi2022long], as well as our own optimised TokenGT model (Table 2). Despite using only half the number of layers as other methods or less, ESA outperformed these baselines and is competitive with the best published models, with further tuning leading to state-of-the-art performance (Table 6).
@@
Learning on 3D atomic systems. We adapt a subset from the Open Catalyst Project (OCP) [zitnick2020introduction, doi:10.1021/acscatal.0c04525] for evaluation, with the full steps in SI 12, including deriving edges from atom positions based on a cutoff, pre-processing specific to crystal systems, and encoding atomic distances using Gaussian basis functions. As a prototype, we compare NSA (MAE of $0.799\pm 0.008$) against Graphormer ($0.839\pm 0.005$), one of the best models that have been used for the Open Catalyst Project [shi2023benchmarking]. These encouraging results on a subset of OCP motivated us to further study modelling 3D atomic systems through the lens of transfer learning.
@@
Transfer learning on frontier orbital energies. We follow the recipe recently outlined for drug discovery and quantum mechanics by [Buterez_Janet_Kiddle_Oglic_Liò_2024] and leverage a recent, refined version of the qm9 HOMO and LUMO energies [Fediai2023] that provides alternative DFT calculations and new calculations at the more accurate GW level of theory. As outlined by [Buterez_Janet_Kiddle_Oglic_Liò_2024], transfer learning can occur transductively or inductively. The transfer learning scenario is important and is detailed in SI 10. We used a 25K/5K/10K train/validation/test split for the high-fidelity GW data, and we trained separate low-fidelity DFT models on the entire dataset (transductive) or with the high-fidelity test set molecules removed (inductive). Since the HOMO and LUMO energies depend to a large extent on the molecular geometry, we used the 3D-aware version of ESA from the previous section and adapted all of our baselines to the 3D setup. Our results are reported in Table 2. Without transfer learning (strategy ‘GW’ in Table 2), ESA and PNA are almost evenly matched, which is already an improvement, since PNA was better for frontier orbital energies without 3D structures (Table 1), while graph transformers perform poorly. When using transfer learning, all methods improve significantly, but ESA outperforms all baselines for both HOMO and LUMO, in both transductive and inductive tasks.
Mixed Graph-level Tasks
@@
Next, we examine a suite of graph-level benchmarks from various domains, including computer vision, bioinformatics, synthetic graphs, and social graphs. Our results are reported in Tables 3 and 8, where ESA generally outperforms all baselines. In the context of state-of-the-art models from online leaderboards, ESA already matches or competes with the best models on well-known datasets such as mnist, malnettiny, and cifar10 without using structural or positional encodings. We also evaluated Exphormer and Polynormer on a representative selection of datasets in SI 4; however, the two baselines proved uncompetitive with ESA.
@@
Table 3: The table reports the Matthews correlation coefficient (MCC) for graph-level classification tasks from various domains, including molecular benchmarks from MoleculeNet (mn) and National Cancer Institute (nci), presented as mean $\pm$ standard deviation over 5 different runs. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g. node/edge features are not integers, which are required for Graphormer and TokenGT). The poor performance of GraphGPS on malnettiny can be explained by our use of one-hot degrees as node features for datasets lacking pre-computed features, while GraphGPS originally used the more informative local degree profile [DBLP:journals/corr/abs-1811-03508]. The highest values and table headers are highlighted in bold. Complete tables, including GCN and GIN, are provided in Tables 7 and 5 with MCC as the metric, and in Tables 8 and 6 with accuracy.
@@
Table 4: The table reports the Matthews correlation coefficient (MCC) for $11$ node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated infected Erdős–Rényi (ER) graphs; details are provided in SI 11). For squirrel and chameleon, we used the filtered datasets introduced by Platonov et al. [platonov2023a] to fix existing data leaks. The highest values and table headers are highlighted in bold. Additional heterophily results are provided in Table 11. A complete table, including GIN and DropGIN, is provided in Table 9 with MCC as the performance metric, and in Table 10 with accuracy.
@@
Table 5: Example of multiple ESA configurations and their impact on performance via three different kinds of benchmarks: a graph-level molecular task (bbbp), a graph-level vision task (mnist), and a node-level heterophilous task (chameleon). In model configurations, M denotes a MAB, S a SAB before/after the PMA module, and P is the PMA module. We illustrate the top performing configurations and their variations, as well as various patterns such as only masked layers or masked/self attention alterations. The performance metric is the Matthews correlation coefficient (MCC). Table headers are highlighted in bold.